OpenStudy (anonymous):

solve the system by the substitution method 15x-10y=5 15y=-4+5x

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

15x-10y=5 15y=-4+5x rearranging the equations 10y=-15x-5 ----(1) 15y=-4+5x -----(2) to Eliminate x,multiply equa(1) by 1 and equa(2) by 3 10y=-15x-5 *1 15y=5x-4 *3 10y=-15x-5 45y=15x-12 add both eqauations 55y=-17 y=-17/55 sub into y into equa(1) x=21/11

OpenStudy (gw2011):

15y=5x-4 y=(5x-4)/15 Substitute this y value into: 15x-10y=5 15x-10[(5x-4)/15]=5 The 10 and 15 can be reduced and you get: 15x-2[(5x-4)/3]=5 15x-[(10x-8)/3]=5 Multiply both sides by 3 to eliminate the denominator 45x-10x+8=15 35x=7 x=7/35=1/5 Substitute this value into the othger equation: 15y=5x-4 15y=5(1/5)-4 15y=1-4 15y=-3 y=-3/15=-1/5 Therefore, x=1/5 and y=-1/5